OpenStudy (anonymous):

write a quadratic equation for x having the given number as the solution. 10 is solution.

OpenStudy (anonymous):

in the form of ax^2+bx+c=0

OpenStudy (anonymous):

x^2-20x+100=0

OpenStudy (anonymous):

ok so far ive gotten x^2-20

OpenStudy (anonymous):

nope wait..

OpenStudy (anonymous):

it has to be ax^2+bx+c=0 form.. its not working i came up with that too and it said it was wrong...idk then

OpenStudy (anonymous):

mine is of that form,ok try this one:2x^2-19x-10=0

OpenStudy (anonymous):

that didnt work but i put in the answer we came up with and said it was wrong and then it gave me the correct asnwer and its the same one we had!! i hate when it does that

OpenStudy (anonymous):

ok now i have one using 7 as the solution, i came up with x^2-16x+64=0. would that be right?

OpenStudy (anonymous):

it would be x^2-16x+63=0

OpenStudy (anonymous):

how did you get 63?

OpenStudy (anonymous):

7^2-(16*7)+63=0 that's the meaning of "7 is a solution of this equation"

OpenStudy (anonymous):

dont you add 8+8 then multiply and that gives you 64?

OpenStudy (anonymous):

that's if u wanted 8 to be a solution of this equation..

OpenStudy (anonymous):

oh lord i typed 7 instead of 8 lol im sorry

OpenStudy (anonymous):

in general there is an infinite number of equations for which 8 is a solution: (x-8)(x+1)=0 (x-8)(x+2)=0 ..... just expand to get your quadratic equation

OpenStudy (anonymous):

got it thanks!